You and your friends have decided to build a skateboard ramp behind your house. You have already purchased $400 in materials and are considering renting an electric saw. It costs $20/hour to rent an electric saw. You estimate it will take 6 hours to build the ramp. If you decide to use your saw, which is a handsaw, you will not have to pay anything to rent an electric saw, but it will take 15 hours to build the ramp. Regardless of the decision, you will need to ask for time off from your job, which pays $12/hour.

Based on the information provided above, answer the following questions.

(a) If you use your handsaw, the explicit costs of the ramp are $ .

(b) The explicit costs of the ramp are $ if you rent an electric saw.

(c)If you use a handsaw, the implicit costs of the ramp are $ .

(d) If you rent an electric saw, the implicit costs of the ramp are $ .

(e) It is economically cheaper to us hand or electric.
Answer #1

(a) Explicit cost ($) = Material cost = 400

(b) Explicit cost ($) = Material cost + Rental cost = 400 + (20 x 6) = 400 + 120 = 520

(c) Implicit cost ($) = Benefit foregone = 15 x 12 = 180

(d) Implicit cost ($) = Benefit foregone = 6 x 12 = 72

(e) Economic cost = Explicit cost + Implicit cost

Economic cost, Hand-saw ($) = 400 + 180 = 580

Economic cost, Electric-saw ($) = 520 + 72 = 592

Since Hand-saw has lower economic cost, Hand-saw is cheaper.